Please help with this problem:
The average of two numbers, a and b is given by A=a+b/2
If the average is 93, and one number is 88, what is the other number?
Thanks,
Janice

Answer by lwsshak3(6527)

The average of two numbers, a and b is given by A=a+b/2
If the average is 93, and one number is 88, what is the other number?
**
let x=other number
(88+x)/2=93
88+x=2*93=186
x=186-88=98